10 points Save Answer A bottle weighed 7 6500 g empty and 99700 g after introduction of an organic liquid with a densay of 0.92 g/mL The balance was equpped with stainless steel weights (dw 80 g ml) Comect the mass of the sample for the effects ofbucyancy Report the mas to the same number of significant figures as the uncorrected mass
Answer #1

Here, the apparent mass of the liquid = 9.9700 – 7.6500 = 2.3200 g.

The same buoyant force acts on the container during both weighing.

Therefore, we need to consider only the force that acts on the 2.3200 g of the liquid.

Now, by substituting, d(air) = 0.0012 g/cm^3

d(obj) = 0.92 g/cm^3

d(wts) = 8.0 g/cm^3

The corrected mass = 2.3200 + 2.3200*(0.0012/0.92 – 0.0012/8.0) = 2.3227 g.